An initial deposit of $9,000 earns 5% interest, compounded monthly. How much will be in the account after 7 years? (Give your answer correct to the nearest cent.)
Answer by rfer(16322)

A=P(1+0.05/12)^nt
A=9000(1.00417)^84
A=9000(1.418315)
A=$12,765.88